# Automatic Intermodal Loading Unit Identification using Computer Vision: A Scoping Review

### v1

@@ 5 Conclusion
This review provides the first synthesis of CV-based ILU identification research by analysing 63 empirical studies published over three decades. There is a clear upward trend for the number of publications with the rise of DL methods and Asia’s dominance as a global trade hub, with 79.71 % of studies originating from the region. Public funding supports over a third of research highlighting a reliance on public investment for transportation related infrastructure. Policymakers could foster progress by funding collaborative initiatives and public-private partnerships, which currently account for just 6.35 % of studies.
Although there are articles published in top journals and conferences, most studies were published in unranked outlets. The main reason for this is the direct application of general purpose CV methods to the ILU identification task. This lack of methodological contribution can be increased by adapting existing methods with modifications that addresses challenges, such as contextless text spotting and real-time processing constraints.
Methodologically, the field has shifted from classical image processing to DL, currently used in more than 90 % of recent studies, thanks to GPU advancements and frameworks like PyTorch. This shift comes with the need of huge datasets to train DL models. Data limitations are the biggest obstacle with 85.71 % of datasets being private and most studies relying on controlled image acquisition setups. The median dataset size is too small for robust DL training, forcing researchers to adapt general text detection and recognition models. Another critical point for improving the quality and credibility of publications is using or developing benchmark datasets. The lack of benchmark dataset makes performance comparisons difficult, leaving gaps in understanding which methods work best in which setting.
The research on ILU identification mainly focuses on fixed camera settings such as OCR-gates. However, recent studies show a shift from fixed to mobile cameras, such as drones or vehicle mounted cameras. This shift can spark more innovation due to new challenges related to dynamic scenes from the technical perspective [84] as well as social perspective such as privacy issues [161, 162, 163]. Camera mounted vehicles offer new subtasks for the ILU identification task such as pose estimation of the ILUs by predicting the real-world coordinates (position) and heading (orientation). This way, a precise monitoring of the terminal is possible.
To conclude, we suggest emphasising standardisation for evaluation with publicly available benchmark datasets, contextless text spotting architectures, real-time processing on edge devices and more focus on vehicle mounted camera setups for a more comprehensive ILU identification and terminal monitoring.
Acknowledgements
The project is supported by the German Federal Ministry for Digital and Transport (BMDV) in the funding program Innovative Hafentechnologien II (IHATEC II) with the funding number 19H23002C.
Disclosure statement
The authors report there are no competing interests to declare.

### v2

@@ 5 Conclusion
This review provides the first systematic synthesis of CV-based ILU identification research by analysing 63 empirical studies published over three decades, and propose a more inclusive terminology. There is a clear upward trend for the quality and number of publications in the recent years driven by the adaptation of CV-based systems, which reduce human error and increase efficiency.
Over the years, the field has shifted from classical DIP to DL thanks to the advancements in computational resources and deep learning frameworks. This shift comes with the need of huge datasets to train DL models and dataset accessibility is the biggest obstacle with 85,71 % of datasets being private. This creates problems for both researchers and practitioners. While large seaports have the means of creating large scale datasets, small sized ports have limitations to create datasets with the same scale, which creates a competitive disadvantage. The median dataset size is too small for robust DL training, forcing researchers to adapt general purpose text detection and recognition models, causing text recognition models to overfit on words with semantic meaning. Limited dataset accessibility also affects the quality and credibility of publications as there are no adapted benchmark datasets. This makes performance comparisons difficult along with the lack of end-to-end analysis, leaving gaps in understanding which methods perform best under which setting. To the best of our knowledge, there are only three publicly available datasets which can be adopted for benchmarking. Considering that two of them were published within 2025, in the near future, the adaptation of these datasets might resolve this problem.
While the rise of DL-based methods brought the need for publicly available datasets, but the shift from static to dynamic image acquisition settings brings other dimensions. The research on ILU identification relied on fixed camera settings such as OCR-gates. However, recent studies are increasingly adopting vehicle mounted camera settings. This shift introduces new challenges like higher variations in object scale caused by changing camera-to-object distances and unknown ILU locations. This is why, other CV and robotics related tasks such as pose estimation (predicting the real-world coordinates and orientation of ILUs) or mission planning for autonomous vehicles (e.g. calculating the most optimal route for identifying ILUs) are now emerging priorities. The seemingly simple problem of identifying ILUs is a fundamental step that has not been taken for autonomous terminal operations. Therefore, addressing these challenges holds transformative potential for transportation and logistics as precise monitoring can optimise the terminal operations, reduce human errors, pave the way for faster turnaround times, lower operational costs and increased resilience against disruptions like misplaces ILUs.